# Dynamics of a Multicomponent Dark Energy Model and the Possibility of EDE like behaviour

### v1

@@ V.5 The Era Shared by Scalar Field and other components
The equilibrium points $ec2_{\pm}$ are defined by $r_{c}=\pm\frac{1}{\sqrt{2}}E_{\pm}D_{\mp}$, $\Omega_{mc}=0$, $\Omega_{rc}=0$, and $\lambda_{c}=A_{\mp}$ and exists for potentials $V_{1}$ and $V_{3}$. There are two distinct types of $\theta_{c}$ at these points: $\theta_{c}=n\pi$ and $\theta_{c}=2n\pi\pm\cos^{-1}\frac{D_{\mp}}{2\sqrt{3}}$. For $\theta_{c}=n\pi$, the eigenvalues of the Jacobian matrix vary depending on whether $n$ is even or odd, but some eigenvalues are always positive, hence the fixed points are unstable. For $\theta_{c}=2n\pi\pm\cos^{-1}\frac{D_{\mp}}{2\sqrt{3}}$, the form of the eigenvalues is complicated and it is not possible to conclude stability analytically for this fixed point for any general choice of the parameters $\alpha$. Thus, stability was assessed for specific values of the parameters $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$ that are considered for the numerical solutions shown in Fig.1 and Fig.3. The Jacobian matrix has at least one positive eigenvalue for our choices of the $\alpha$ parameters, making $ec2_{\pm}$ unstable. These points also correspond to an accelerating universe.
VI EDE like behavior of the scalar field
@@
In our analysis of the phase space for the current model, which includes a nonminimally coupled scalar field and the cosmological constant as potential candidates for dark energy, we observed that there is no late-time attractor completely dominated by the scalar field. The only possible late-time attractors are either completely dominated by the cosmological constant or a scenario where there is only the cosmological constant and scalar field present in the universe, the rest of the components are absent. Although there are epochs where the scalar field and other components coexist, these correspond to the fixed points, which are saddle or unstable.
In the following, we aim to explore the possibility of the scalar field contributing to the universe’s energy budget prior to the onset of the dark energy dominated epoch. This is motivated by the presence of fixed points in the system that corresponds to the eras that might be potentially shared between other components and the scalar field, suggesting an early dark energy (EDE) like behaviour of the scalar field.
@@
The evolution of the density parameters for various components of the universe corresponding to all the classes of potentials are shown in Fig.4 to Fig.11 respectively. For the classes of potentials, $V_{1}$ $V_{6}$ we can see the EDE like behaviour of the scalar field appears naturally in the deep-matter dominated era, which is almost $10\%$ of the total energy budget of the universe. But, at present the contribution of the scalar field is negligible. Whereas for the classes of potentials, $V_{7}$ and $V_{8}$ also the similar behaviour of the scalar field appears naturally in the deep matter-dominated era, which is almost $10\%$ of the total energy budget of the universe. Contrary to the previous case for these two classes of potentials at present as well as near future the scalar field has a non-negligible contribution to the total energy budget of the universe. The redshift parameter values for matter-radiation equality and the occurrence of EDE-like behavior of the scalar field are provided in Table-6 for the numerical solutions obtained for all classes of potentials.
@@
We have also plotted the evolution of different cosmological parameters like the deceleration parameter in Fig.12, the equation of state of the scalar field in Fig.13 and the total equation of state in Fig.14 for all potentials. Those plots are drawn for the same values of the parameters $\alpha$ for which Fig.4 to Fig.11 are drawn. For the evolution of the deceleration parameter, one can see that its value has become negative in the recent past from positive, showing the accelerating expansion of the universe. It can be seen that the transition of the universe from the decelerated to accelerated phase remains smooth throughout except for the epoch where the scalar field shows EDE like behaviour. During this epoch, the deceleration parameter shows some increments suggesting more deceleration of the universe due to the activation of the scalar field. Later, it started to decrease and become negative with time. From the plot of the EOS of the scalar field in Fig.13 one can see that in the very early universe, the scalar field behaves like a stiff fluid with EOS $w_{\phi}=1$ then it becomes $w_{\phi}=-1$. During the EDE like behaviour of the scalar field, it increases and becomes positive and the scalar field does not contribute to the acceleration of the universe since $w_{\phi}>-1/3$ rather it was slowing down the expansion of the universe. This can also be seen from the plot of the $w_{tot}$ in Fig.14 since during this epoch the $w_{tot}$ becomes positive and starts to decrease again.
In Fig.15 we have shown the evolution of the $H(z)$ vs $z$ for all the classes of potentials we have considered here. It can be seen that these potentials fit the data quite well, at least during the late time evolution. We acknowledge that a comprehensive parameter estimation using current cosmological data is necessary to constrain the parameters and comment more accurately on this EDE like behaviour of the scalar field, which is beyond the scope of the current work. This will be addressed in future research.
@@ VII Conclusion
In this study, we explore the dynamics and phase-space characteristics of a multicomponent dark energy model. This model features a dark sector that includes a minimally coupled canonical scalar field and a cosmological constant. The equations are transformed into autonomous systems and subsequently into a polar form. We have considered a parameterization for the potentials of the scalar field that helps us to study a broad spectrum of potentials in a single setup.
@@
We also investigated numerically the possibility of EDE like behaviour for all the different classes of potentials considered in this work and observed that this EDE like behaviour of the scalar field occurs deep within the matter-dominated era, not near the recombination period. We also found that during this EDE like era of the scalar field, it behaves more like matter rather than dark energy. During this era, the universe becomes more decelerated compared to the standard case. The total EOS of the universe also shows a transition from negative to positive value during this epoch. A detailed examination of this EDE like behaviour of the scalar field will require a comprehensive cosmological data analysis with the recent data, which we will present in the future.
Figure 4: Plot of the density parameters for different components of universe for the potential $V_{1}$.
@@
Figure 12: Plots of the deceleration parameter for early dark energy with respect to different classes of potentials.

### v3

@@ V.5 The Era Shared by Scalar Field and other components
The equilibrium points ‘$ec2_{\pm}$’ are defined by $r_{c}=\pm\frac{1}{\sqrt{2}}E_{\pm}D_{\mp}$, $\Omega_{mc}=0$, $\Omega_{rc}=0$, and $\lambda_{c}=A_{\mp}$ and exists for potentials $V_{1}$ and $V_{3}$. There are two distinct types of $\theta_{c}$ at these points: $\theta_{c}=n\pi$ and $\theta_{c}=2n\pi\pm\cos^{-1}\frac{D_{\mp}}{2\sqrt{3}}$. For $\theta_{c}=n\pi$, the eigenvalues of the Jacobian matrix vary depending on whether $n$ is even or odd, but some eigenvalues are always positive, hence the fixed points are unstable. For $\theta_{c}=2n\pi\pm\cos^{-1}\frac{D_{\mp}}{2\sqrt{3}}$, the form of the eigenvalues is complicated, and it is not possible to conclude stability analytically for this fixed point for any general choice of the parameters $\alpha$. Thus, stability was assessed for specific values of the parameters $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$ that are considered for the numerical solutions shown in Fig.1 and Fig.3. The Jacobian matrix has at least one positive eigenvalue for our choices of the $\alpha$ parameters, making ‘$ec2_{\pm}$’ unstable. These points also correspond to an accelerating universe.
VI Numerical investigation
@@
In our analysis of the phase space for the current model, we observed that there is no late-time attractor completely dominated by the scalar field. The only possible late-time attractors are either completely dominated by the cosmological constant or a scenario where there is only the cosmological constant and scalar field present in the universe; the rest of the components are absent. Although there might be some epochs where the scalar field and other components coexist, represented by saddle or unstable fixed points. For example, the fixed points in Table:2 where $r_{c}$ and some other components like $\Omega_{mc}$ or $\Omega_{rc}$ are non-zero. In the following, we have numerically evolved the system to investigate the evolution of the different cosmological parameters, and we will see, as expected from the fixed point analysis, that there are possibilities of some eras of the universe which are shared by the scalar field and other components of the universe much before the standard era of dark energy domination.
@@
The evolution of the density parameters for various components of the universe corresponding to all the classes of potentials are shown in Fig.4 to Fig.11 respectively. For the classes of potentials, $V_{1}$ $V_{6}$ one can see the scalar field contribute to the energy budget of the universe earlier than the expected dark energy domination epoch for a short period resembling EDE-like behaviour of the scalar field appearing naturally in the deep-matter dominated era, which is almost $10\%$ of the total energy budget of the universe. But, at present the contribution of the scalar field is negligible for these potentials. Whereas for the classes of potentials, $V_{7}$ and $V_{8}$ also the similar behaviour of the scalar field appears naturally in the deep matter-dominated era, which is almost $10\%$ of the total energy budget of the universe. Contrary to the previous case, for these two classes of potentials, at present as well as near future, the scalar field has a non-negligible contribution to the total energy budget of the universe. The redshift parameter values for matter-radiation equality and the occurrence of this behavior of the scalar field are provided in Table-6 for all classes of potentials.
@@
We have also plotted the evolution of different cosmological parameters like the deceleration parameter in Fig.12, the equation of state of the scalar field in Fig.13 and the total equation of state in Fig.14 for all potentials. Those plots are drawn for the same values of the parameters $\alpha$ for which Fig.4 to Fig.11 are drawn. For the evolution of the deceleration parameter, one can see that its value has become negative in the recent past from positive, showing the accelerating expansion of the universe. It can be seen that the transition of the universe from the decelerated to the accelerated phase remains smooth throughout except for the epoch where the scalar field shows early time contribution. During this epoch, the deceleration parameter shows some increments suggesting more deceleration of the universe due to the activation of the scalar field. Later, it started to decrease and become negative with time. From the plot of the EOS of the scalar field in Fig.13 one can see that in the very early universe, the scalar field behaves like a stiff fluid with EOS $w_{\phi}=1$ then it becomes $w_{\phi}=-1$. During the early contribution period of the scalar field, it increases and becomes positive and the scalar field does not contribute to the acceleration of the universe since $w_{\phi}>-1/3$ rather it was slowing down the expansion of the universe. This can also be seen from the plot of the $w_{tot}$ in Fig.14 since during this epoch the $w_{tot}$ becomes positive and starts to decrease again.
In Fig.15 we have shown the evolution of the $H(z)$ vs $z$ for all the classes of potentials together with the observed data from different observations for comparison. Please see Section:B for more details about the data used in this plot. It can be seen that even though the scalar field contributed during the matter-dominated era, these potentials fit the data quite well, at least during the late time evolution.
A similar approach to study the early dark energy behavior of the scalar field using dynamical system analysis was used in [76, 77]. Both works show that EDE-like behavior of the scalar field before the recombination era can be achieved, but it is subjected to a high degree of approximations. One of the advantages of our approach is that one can find EDE-like behaviour of the scalar field more generically without any approximation and the result applies to a wide class of potentials.
@@ VII Conclusion
In this study, we explore the dynamics and phase-space characteristics of a multi-component dark energy model. This model features a dark sector that includes a minimally coupled canonical scalar field and a cosmological constant. The equations are transformed into autonomous systems and subsequently into polar form. Since we are interested in making our analysis valid for a wide class of potentials, enabling us to study the multi-component models from a more general point of view, we have considered a parameterization for the potentials of the scalar field that helps us study a broad spectrum of potentials in a single setup.
@@
We also investigated numerically the evolution of different cosmological parameters and compared it with the observed data. Most interestingly, our findings show the early time contribution of the scalar field occurring deep within the matter-dominated era, not near the recombination period. We also found that during this era of the scalar field, it behaves more like matter rather than dark energy, and the universe becomes more decelerated compared to the standard case. The total EOS of the universe also shows a transition from negative to positive values during this epoch. We acknowledge that a comprehensive parameter estimation for the choice of $\alpha$ parameters using a Boltzmann code like CLASS and current cosmological data is necessary to comment more accurately on this early-time contribution of the scalar field and how it could affect the overall evaluation and structure formation of the universe. Since this study focuses on the investigation of the phase space behaviour of the model, it is beyond the scope of the current work. It will be addressed in future research.
Figure 4: Plot of the density parameters for different components of the universe for the potential $V_{1}$.
@@
Figure 12: Plots of the deceleration parameter with respect to different classes of potentials.